• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name E. Schulte

    Log entry time 16:01:15 on November 01, 2008

    Entry number 245962

    keyword=Day Shift Summary for November 1, 2008

    08:00 Shift starts with smooth running on hydrogen reference cell. At
    start of shift there was approximately 1.5 hours remaining in this
    configuration. Beam current is at 0.9 uA and a little trippy.

    09:25 We are continuing to take data on the reference cell. We are
    waiting for the RC because we have some issues with the run plan.
    Collimators need to be removed/replaced/moved around before the next step
    can be performed. There is some confusion about what is needed to
    accomplish this task. At some point today, RadCon needs to be called in
    to facilitate this. RC has been called, and is supposed to be coming in
    soon...So we will patiently await his advice.

    09:30 The beam trip frequency seems to be increasing over the last few

    09:37 Beam trips seem to have settled down.

    09:44 The RC's here! Called for a Controlled Access to insert the target

    10:00 The ARM is here. They will have an escorted access this time.

    10:20 Found that Compton slot 15 channel 10 was tripped off. This was
    missed because the Alarm Handler had been closed. Vince found the
    problem. The Alarm Handler is now back on and the channel was reset.

    10:23 Vince starts making HV changes...Hopefully he'll remember to put
    them in the logbook.

    11:24 Joe is out of the hall & we're changing to BeO target.

    11:25 Requested Beam Permit from MCC.

    11:32 Requested 1uA with raster off.

    11:41 After a quick check of the beam position on the BeO target we
    requested a change of target to the H2 reference cell.

    11:45 Requested 1uA with 4x4 raster on H2 reference cell from MCC. We
    will take 1.5 hours of data on resumption of beam delivery.

    11:50 BigBite wire chambers tripped.

    12:15 BigBite wire chamber tripped again. Fault was cleared. Then it
    tripped again and we're clearing the faults. Called MCC to as them to
    please ramp the beam back more slowly to try to avoid the BigBite wire
    chamber faults.

    12:22 Once all those were reset, L14.7 tripped. It was reset. MCC seems
    to be having some problems. They've been recovering from a trip for a
    while now.

    12:30 BigBite is not a happy camper. L14.7 & L15.3 both tripped off
    again. They are being reset. On a happier note, the beam is back!

    12:37 MCC called to let us know that they are working on the beam ramp
    rate (to slow it down) for the BigBite wire chamber trips. Apparently,
    the ramp rate was turned off because we were only taking 1uA.

    12:49 BigBite is really not happy. A few wire chamber channels have been
    tripping since we came out of the Controlled Access. We have requested
    that MCC take the beam away and an expert will be called.

    13:02 Requested 1uA beam back from MCC. Beam came back at 1uA without a
    peep from the BigBite wire chambers. At least thus far. We will run
    this for about 1.5 hours.

    13:15 Well, we survived one beam trip, but now the beam has become trippy
    again, it looks like.

    13:34 Beam's having trouble again. It's been about 10 minutes beam free.

    13:52 Called MCC to find out what happened to the beam. They are having
    a problem with an RF cavity. They gave no estimate on when the beam will
    be back.

    14:03 Alexandre asked if we could take a RICH pedestal run. We will
    return to normal when the beam comes back.

    14:32 Still no beam.

    14:33 Beam's back! Return DAQ to TwoArm_FB configuration and resume
    taking data. So far, BigBite has been reasonably happy.

    14:47 BigBite wire chamber channel L14.7 tripped again taking L14.6 with
    it on reset. L14.2 & L14.3 also required a reset. They're coming back
    up. This was clearly not associated with any obvious beam misstearing.
    There was a trip in the recent past, but it wasn't clearly associated
    with that either, as far as we can tell.

    14:53 BigBite wire chamber channel L14.6 tripped again taking L14.2 with
    it. This one did seem to be correlated with a beam trip.

    14:55 CODA end of run failed!

    14:56 BigBite wire chamber channels L14.2, .3, .6, .7 all tripped at once
    this time. We are also having trippy beam, and CODA stopped working.

    15:04 Joe reported that we're at 50% polarization as of the last NMR

    15:07 MCC needs to tune beam into Hall B, so we will loose the beam for a

    15:15 MCC called to let us know that they will be sending 1uA.
    BigBite wire chamber channels L14.2,.3,.6.,7 tripped again on return of
    beam delivery (it seems). L15.0, .2, .3 also tripped off this time.

    15:27 After 2 resets we're ready to try taking data again. The beam
    looks "relatively" stable at 0.9uA.

    15:37 Channels L14.2, .3, .6, .7 and L15.2 & .3 tripped again. It's
    taken 2 resets so far to get them back (and they're still working on it).

    15:50 Resume taking data. And then the beam tripped.

    16:00 We've been through several "trip" cycles & the wire chambers
    haven't tripped off. Not sure what to make of that. We have about 2M
    events on the reference cell (with H2), maybe 40 minutes or so? The beam
    has been quite poor all shift. Between the bad beam and the cantankerous
    wire chambers, we haven't gotten much dun since the Controlled Access
    earlier this shift.